OpenStudy (anonymous):

$400 is divided b/w Kas, Jas and Jae so that Kas has twice as much as Jas and Jas has 3 times as much as Jae. how much does Jas receive?

OpenStudy (loser66):

what does b/w mean?

OpenStudy (anonymous):

between

OpenStudy (loser66):

 got this part?

OpenStudy (loser66):

try to get it, if you cannot, I cannot go further

OpenStudy (anonymous):

its 6x:x:3x so 400/10 = 40 thts jaes share idk hw to find the rest

OpenStudy (loser66):

yes, but it's not 6x:x:3x it is 6x+3x+x =10x = 400 , so x = 400/10 = 40 just notation. and you are right. we have Jae is 1x = 40 Jas is 3x = 3*40 = 120 Kas is 6x = 6*40 = 240 ----------------------------------- total 400 \(\surd \)

OpenStudy (anonymous):

ohhh okay thanks :)

OpenStudy (loser66):

np